# Understanding Stellar Mass-Metallicity and Size Relations in Simulated Ultra-Faint Dwarf Galaxies

### v1

@@ 3.2.5 MDF with different SN energy
Figure 13: Top panel: Surface brightness map of the UFD analog, with star particles overlaid and color-coded by their progenitor halo. Artificial background stars for fitting are excluded from this plot. Bottom panel: The surface brightness density profile, which is calculated by binning the elliptical radius from the center of the simulated UFD analog. Data including background stars (large black circles) and excluding background stars (small grey circles) are shown. The solid red line represents the fit, including background stars, while the dashed line represents a simple exponential fit, $\propto\exp(-r/r_{\rm h})$ with $r_{\rm h}$ from the fit.
@@
UFD $M_{\star}$ $\rm\langle[Fe/H]\rangle$ $r_{\rm h,dir}$ $r_{\rm h,fit}$ $\epsilon_{\rm h,fit}$
analog $[10^{4}\rm~{}\mbox{M}_{\odot}]$ [pc] [pc]
Halo1 0.24 -2.19 89 $87\pm 16$ $0.34\pm 0.16$
Halo2 1.22 -2.78 228 $182\pm 17$ $0.44\pm 0.06$
Halo3 0.52 -2.98 175 $251\pm 33$ $0.58\pm 0.06$
Halo4 1.58 -2.65 176 $180\pm 13$ $0.23\pm 0.07$
Halo5 2.14 -2.71 1115 $754\pm 201$ $0.57\pm 0.11$
Table 2: Structural parameters of the simulated UFDs at $z=0$. The columns represent the following parameters: (1) $M_{\star}$: Stellar mass. (2) $\rm\langle[Fe/H]\rangle$: Average stellar iron-to-hydrogen ratio for all stars within $r_{\rm vir}$. (3) $r_{\rm h,\ dir}$: 2D half-light radius based on the direct method (4) $r_{\rm h,\ fit}$: 2D half-light radius based on the fitting method. (5) $\epsilon_{\rm h,\ fit}$: Ellipticity.
@@ 3.3.1 Half-light radius
In addition to MZR, both observations and simulations of UFDs have shown discrepancies in the size-luminosity relation, with cosmological simulations struggling to reproduce the observed compactness ($r_{h}\lesssim 50\rm\ pc$). Simulated UFDs typically exhibit larger sizes, spanning several hundred parsecs (e.g., Jeon et al., 2017; Prgomet et al., 2022; Sanati et al., 2023), and in some cases, sizes approach nearly 1 kpc (e.g., Wheeler et al., 2019; Jeon et al., 2021b). A possible cause for this size discrepancy may be differences in the methods used to derive the half-light radius. For simulated UFDs, the half-light radius is typically measured by finding the radius where the stellar mass within the virial radius is halved while applying a constant mass-to-light ratio to the stellar particles. On the other hand, observed UFDs are modeled using specific functional forms, such as an exponential profile, and maximum likelihood fits are performed on the probability density functions with free parameters to describe their spatial structures. To address this methodological difference between simulation and observation, we apply the same method commonly used in observations to derive the structural parameters of stellar distributions, utilizing maximum likelihood estimation techniques. For a detailed mathematical explanation, we refer readers to Martin et al. (2008).
@@
where $r$ is the elliptical radius and $r_{\rm e}$ is the exponential scale radius, related to the half-light radius by $r_{\rm h}=1.68r_{\rm e}$. The parameters $\Sigma_{0}$ and $\Sigma_{\rm b}$ represent the central and background stellar densities, respectively, with $\Sigma_{0}$ given by $\Sigma_{0}=N_{\star}/(2\pi r_{\rm e}^{2}(1-\epsilon))$. Here, $N_{\star}$ is the total number of stars and $\epsilon$ is the ellipticity, defined as $\epsilon=1-b/a$, with $b/a$ being the minor-to-major axis ratio of the galaxy. The elliptical radius is calculated as
@@
We should highlight the importance of including artificial background stars and their impact on fitting results. Artificial background stars [^4] To avoid confusion among observers, we will refer to ”artificial background stars,” used solely for simulated UFDs, simply as ”background stars.” refer to stars added to simulated galaxies to replicate the observational conditions of UFDs. Without accounting for background stars, we find that when there are extended stars in UFDs, maximum likelihood estimation produces an improper fit due to the lack of a suitable exponential stellar density profile that adequately captures both the inner, compact distribution and the outer, extended structure. Consequently, the estimated half-light radius is significantly larger than observed values typically seen in UFDs.
@@
For instance, Halo1 contains seven extremely extended star particles distributed over a few kiloparsecs. Including these stars and fitting only the star particles within the virial radius of the simulated galaxies, without accounting for background stars, results in a half-light radius of $r_{\rm h,\ fit}=508\ \rm pc$, making it difficult to accurately capture both the inner and outer profiles simultaneously. This estimated half-light radius is excessively larger than that obtained via the direct method, which yields a much smaller radius of $r_{\rm h,\ dir}=86\ \rm pc$. In actual observations, these seven extended stars would likely be misidentified as background stars due to their low surface brightness and thus not included as members of the galaxy. To take an approach similar to observations, we artificially generate background stars using Poisson random sampling, assuming a background stellar density of $\rm 0.1\ arcmin^{-2}$, as an initial value ($\Sigma_{\rm b,0}$) during the fitting process. The value is arbitrary but representative of the background density of observed dwarf galaxies such as Draco (e.g., Ségall et al., 2007). This approach reduces the half-light radius when applying the fitting method, as the extended stars are less likely to be included as members during the fitting.
The estimated structural parameters of the simulated UFD analogs are summarized in Table 2. We compare the half-light radii obtained using the direct method, $r_{\rm h,\ dir}$, with those, $r_{\rm h,\ fit}$, derived by fitting the best-fitting parameters from equation (3). For Halo1 and Halo4, both methods yield nearly identical half-light radii. However, the other three halos exhibit discrepancies between $r_{\rm h,\ dir}$ and $r_{\rm h,\ fit}$. For example, the half-light radii from the direct method for Halo2 and Halo5 are 25% and 47% larger, respectively, than those derived from the fitting method. These differences are primarily due to the higher fraction of extended stars, which are more likely to be identified as background stars in the fitting method, resulting in a smaller estimated radius.
UFD analog $r_{\rm h}[\rm pc]$ $r_{s}[\rm pc]$ $r_{\rm trans}[\rm pc]$ $B$ Halo1 $69\pm 5$ $482\pm 48$ 287 0.002 Halo2 $122\pm 37$ $690\pm 46$ 361 0.012 Halo3 $115\pm 8$ $365\pm 13$ 319 0.023 Halo4 $104\pm 13$ $347\pm 23$ 322 0.014 Halo5 $146\pm 13$ $747\pm 38$ 312 0.042
@@
Figure 14: The representative surface brightness profile of Halo4 is modeled using a two-component approach. The actual density profile is shown as black-filled circles at binned radii. The inner component is depicted in red, while the outer component, characterized by a scale radius of $r_{\rm s}=346$ pc, is illustrated in blue. The transition between the inner and outer components occurs at a radius of $r_{\rm trans}=322$ pc.
@@
In the top panels of Fig. 13, we illustrate the spatial distribution of our UFD analogs, with star particles from different progenitor halos represented in distinct colors overlaid on the surface brightness maps. Background stars used for fitting are excluded from this plot. Among the simulated UFD analogs, Halo1 exhibits the most compact features, with a half-light radius of $r_{\rm h,\ fit}\approx 87$ pc and the smallest ellipticity of $\epsilon_{\rm fit}=0.34$, indicating a more spherical distribution of stellar particles. Conversely, stars in Halo5 are the most extended, with $r_{\rm h,\ fit}\approx 754$ pc, and tend to deviate from a spherical shape, displaying a more elongated distribution with ellipticity reaching $\epsilon_{\rm fit}=0.57$. Note that Halo6, having a similar halo mass to Halo5, is excluded from the size relation analysis to avoid redundancy.
The bottom panels of Fig. 13 display the surface brightness profile as a function of elliptical radius from the center of the simulated UFD analogs. The black large circles (including background stars) and grey small circles (excluding background stars) represent the surface brightness directly computed from stars in each UFD analog. The solid red line depicts the derived exponential fit using maximum likelihood estimation, including background stars, while the red dashed line indicates a simple exponential fit, $\propto\exp(-r/r_{h})$, without background stars. We verify that accurately estimating the spatial structures of simulated UFDs using observational methodologies necessitates the inclusion of background stars to achieve a proper fit. Furthermore, it is evident that, due to extended stars at large radii (indicated by grey circles), a simple extension of the inner profile (red dashed line) no longer accurately fits the total stellar density profile, especially the outer structures. Therefore, a different density profile that simultaneously represents both the inner, compact structure and the outer, extended structure would be more appropriate (e.g., Kang & Ricotti, 2019; Ricotti et al., 2022; Jensen et al., 2024).
@@ 3.3.3 Size-luminosity relation
We compare the half-light radii of our simulated UFD analogs with both observational data and other theoretical studies. The observational results include an updated catalog of dwarf galaxies from McConnachie (2012), a compilation of dwarf galaxies by Simon (2019), and observed UFDs reported by Richstein et al. (2022, 2024). While Richstein et al. (2022, 2024) use azimuthally averaged half-light radii, we adopt the elliptical half-light radius, $a_{\rm h}$, from their work to maintain consistency with other works using this definition. The simulated galaxies are represented in different colors (Jeon et al., 2017, 2021b; Wheeler et al., 2019; Applebaum et al., 2021; Prgomet et al., 2022; Sanati et al., 2023). For the V-band luminosity, we uniformly apply a mass-to-light ratio of 2 to our results and the other simulated galaxies.
@@
As illustrated in Fig. 17, we compute two half-light radii for each UFD analog using the direct method ($r_{\rm h,\ dir}$, denoted as circles) and the two-component fitting method ($r_{\rm h}$, denoted as crosses). While the $r_{\rm h}$ values are consistently smaller than the $r_{\rm h,\ dir}$ values for all the simulated UFDs, the former closely match the observational data. This consistency suggests that the large half-light radii reported in other simulations (e.g., Wheeler et al., 2019; Sanati et al., 2023) could also be reduced by applying the two-component fitting method if extended structures are present. Although our derived values using the two-component profile successfully match the observational data at the upper range, the average values remain larger ($r_{h}\sim$ a few $100\ \rm pc$) compared to the typical observed values ($r_{h}\lesssim 100\ \rm pc$). These larger sizes in the UFD regime predicted by theoretical work imply that observed UFDs may possess more extended yet unresolved structures. This possibility is further supported by recent observational studies revealing the presence of extended structures in UFDs (e.g., Tau et al., 2024; Jensen et al., 2024).
@@
In particular, our simulations still struggle to reproduce the highly compact sizes with $r_{\rm h}\lesssim 50$ pc. This discrepancy is particularly notable for UFDs with $M_{\star}\sim 10^{3}~{}\mbox{M}_{\odot}$, where observational studies consistently suggest sizes around $r_{\rm h}\sim 30$ pc. Numerical heating caused by different mass resolutions between DM and star particles could contribute to the larger sizes (e.g., Revaz & Jablonka, 2018; Binney & Knebe, 2002). Ludlow et al. (2019b) showed that the heavier DM particles tend to sink to the center of simulated galaxies, causing the less massive stellar particles to gradually diffuse outward. This process results in a systematic increase in the half-light radius over time until it reaches the convergence radius, $r_{\rm conv}\approx 0.055\times l(z)$ (Ludlow et al., 2019a), where $l(z)$ represents the mean interparticle spacing, defined as $l=L_{\rm box,\ phy}/N_{\rm part}^{1/3}$. Here, $L_{\rm box,\ phy}$ is the physical size of the simulation box, and $N_{\rm part}$ is the number of particles. In our study, the zoom-in region, with a resolution of a $2048^{3}$ grid, yields a convergence radius of $r_{\rm covn}\approx 59$ pc, including baryon particles. Given that our estimated $r_{\rm h}$ values are larger than this limit, the impact of numerical heating may not be significant in our work.
@@
We also mention that Jeon et al., 2021b, a, where a simulation setup similar to ours ($m_{\rm DM}=500~{}\mbox{M}_{\odot}$ and $m_{\rm gas}=63~{}\mbox{M}_{\odot}$), demonstrated the existence of simulated UFDs with smaller sizes ($r_{\rm h}<50$ pc). In these studies, numerical heating caused by differing mass ratios between DM and star particles was not a concern, as each star particle was modeled as a SSP with an initial mass of $m_{\star}\approx 500~{}\mbox{M}_{\odot}$, formed by accreting gas particles, giving rise to the same mass as a gas particle. However, the total stellar masses in these UFD analogs consisted of fewer than 10 particles for a stellar mass of $M_{\star}\lesssim 2\times 10^{3}~{}\mbox{M}_{\odot}$. This small number of particles representing the total stellar mass may contribute to the lower $r_{\rm h}$ values.
@@
Figure 18: Stellar metallicity as a function of 3D radial distance from the center at $z=0$. Here, the metallicity gradient is not observed in any simulated UFDs; rather, these galaxies exhibit a wide range of metallicity distributions, reflecting the diverse origins of stars from different progenitor halos. Specifically, in Halo5, stars originating from prog5 (indicated in purple) are found dispersed across vast regions extending from 1 to 10 kpc.

### v2

@@ 3.2.5 MDF with different SN energy
Figure 13: Top panel: Surface brightness map of the UFD analog, with star particles overlaid and color-coded by their progenitor halo. Artificial background stars used for fitting are excluded from this plot. Bottom panel: The surface brightness density profile, calculated by binning the elliptical radius from the center of the simulated UFD analog. Data are shown for both cases, including background stars (red circles) and excluding background stars (grey circles). The solid red line represents the fit with the background, while the grey line represents the fit without the background. The latter one, which considers all stars within the virial radius of the UFD analog as the member stars, results in a poor fit with an exceptionally large radius.
@@
UFD $M_{\star}$ $\rm\langle[Fe/H]\rangle$ $r_{\rm h,dir}$ $r_{\rm h,fit\ (w/)}$ $\epsilon_{\rm h,fit\ (w/)}$ $r_{\rm h,fit\ (w/o)}$ $\epsilon_{\rm h,fit\ (w/o)}$
analog $[10^{4}\rm~{}\mbox{M}_{\odot}]$ [pc] [pc] [pc]
Halo1 0.24 -2.19 89 $87\pm 16$ $0.34\pm 0.16$ $516\pm 71$ $0.25\pm 0.12$
Halo2 1.22 -2.78 228 $182\pm 17$ $0.44\pm 0.06$ $667\pm 39.5$ $0.31\pm 0.05$
Halo3 0.52 -2.98 175 $251\pm 33$ $0.58\pm 0.06$ $390\pm 36$ $0.53\pm 0.05$
Halo4 1.58 -2.65 176 $180\pm 13$ $0.23\pm 0.07$ $445\pm 23.5$ $0.40\pm 0.04$
Halo5 2.14 -2.71 1115 $754\pm 201$ $0.57\pm 0.11$ $1926\pm 86$ $0.62\pm 0.02$
Table 2: Structural parameters of the simulated UFDs at $z=0$. The columns represent the following parameters: (1) UFD analog name. (2) $M_{\star}$: Stellar mass. (3) $\rm\langle[Fe/H]\rangle$: Average stellar iron-to-hydrogen ratio for all stars within $r_{\rm vir}$. (4) $r_{\rm h,\ dir}$: 2D half-light radius based on the direct method. (5)–(8): Results from the fitting method: half-light radius and ellipticity derived with background stars ($r_{\rm h,\ fit\ (w/)}$, $\epsilon_{\rm h,\ fit\ (w/)}$) and without background stars ($r_{\rm h,\ fit\ (w/o)}$, $\epsilon_{\rm h,\ fit\ (w/o)}$).
@@ 3.3.1 Half-light radius
In addition to MZR, both observations and simulations of UFDs have shown discrepancies in the size-luminosity relation, with cosmological simulations struggling to reproduce the observed compactness ($r_{\rm h}\lesssim 50\rm\ pc$). Simulated UFDs typically exhibit larger sizes, spanning several hundred parsecs (e.g., Jeon et al., 2017; Prgomet et al., 2022; Sanati et al., 2023), and in some cases, sizes approach nearly 1 kpc (e.g., Wheeler et al., 2019; Jeon et al., 2021b). A possible cause for this size discrepancy may be differences in the methods used to derive the half-light radius. For simulated UFDs, the half-light radius is directly measured by finding the radius where the stellar mass within the virial radius is halved while applying a constant mass-to-light ratio to the stellar particles. On the other hand, observed UFDs are modeled using specific functional forms, such as an exponential profile, and maximum likelihood fits are performed on the probability density functions with free parameters to describe their spatial structures. For convenience, we refer to the former as the direct method and the latter as the fitting method. To address this methodological difference between simulation and observation, we apply the same method commonly used in observations to derive the structural parameters of stellar distributions, utilizing maximum likelihood estimation techniques. For a detailed mathematical explanation, we refer readers to Martin et al. (2008).
@@
where $r$ is the elliptical radius and $r_{\rm e}$ is the exponential scale radius, related to the half-light radius by $r_{\rm h}=1.68\ r_{\rm e}$. The parameters $\Sigma_{0}$ and $\Sigma_{\rm b}$ represent the central and background stellar densities, respectively, with $\Sigma_{0}$ given by $\Sigma_{0}=N_{\star}/(2\pi r_{\rm e}^{2}(1-\epsilon))$. Here, $N_{\star}$ is the total number of galaxy member stars, and $\epsilon$ is the ellipticity, defined as $\epsilon=1-b/a$, with $b/a$ being the minor-to-major axis ratio of the galaxy. The elliptical radius is calculated as
@@
Here, we should highlight the importance of including artificial background stars and their impact on fitting results. Artificial background stars [^5] To avoid confusion among observers, we will refer to ”artificial background stars,” used exclusively in simulated UFDs, simply as ”background stars”. We also note that the term ”background stars” in this paper encompasses both background and foreground stars in real observations of UFDs. refer to stars added to simulated galaxies to replicate the observational conditions of UFDs. Without accounting for background stars, we find that when there are extended stars in UFDs, maximum likelihood estimation produces an improper fit due to the lack of a suitable exponential stellar density profile that adequately captures both the inner, compact distribution and the outer, extended structure. Consequently, the estimated half-light radius is significantly larger than observed values typically seen in UFDs.
@@
For instance, Halo1 contains seven extremely extended star particles distributed over a few kiloparsecs. Including these stars and fitting only the star particles within the virial radius of the simulated galaxies, without accounting for background stars, results in a half-light radius of $r_{\rm h,\ fit}=516\ \rm pc$, making it difficult to accurately capture both the inner and outer profiles simultaneously. This estimated half-light radius is excessively larger than that obtained via the direct method, which yields a much smaller radius of $r_{\rm h,\ dir}=89\ \rm pc$. In actual observations, these seven extended stars would likely be misidentified as background stars due to their low surface brightness and thus not included as members of the galaxy. To take an approach similar to observations, we artificially generate background stars using Poisson random sampling, assuming a background stellar density of $\rm 0.1\ arcmin^{-2}$, as an initial value ($\Sigma_{\rm b,0}$) during the fitting process. The value is arbitrary but representative of the background density of observed dwarf galaxies such as Draco (e.g., Ségall et al., 2007). This approach reduces the half-light radius when applying the fitting method, as the extended stars are less likely to be included as members during the fitting.
The estimated structural parameters of the simulated UFD analogs are summarized in Table 2. We compare the half-light radii derived by fitting the best-fitting parameters from equation (3), both with and without background stars. Additionally, we calculated $r_{\rm h,\ dir}$ to assess the extent to which the fitting results deviate from the directly calculated radius. The comparison between $r_{\rm h,fit\ (w/o)}$ and $r_{\rm h,fit\ (w/)}$ clearly shows that excluding background stars in the fit leads to a systematically higher value, on average, three times larger. The most significant difference is observed in Halo1, where $r_{\rm h,fit\ (w/)}=516$ pc is approximately six times larger than $r_{\rm h,fit\ (w/o)}=87$ pc. Comparing $r_{\rm h,fit\ (w/)}$ and $r_{\rm h,\ dir}$, we find similar values, with any deviations likely arising from the actual distribution not perfectly following an exponential profile.
UFD analog $r_{\rm h}[\rm pc]$ $r_{s}[\rm pc]$ $r_{\rm trans}[\rm pc]$ $B$ Halo1 $84\pm 9$ $964\pm 76$ 391 0.001 Halo2 $117\pm 4$ $1302\pm 38$ 467 0.008 Halo3 $109\pm 8$ $587\pm 26$ 283 0.028 Halo4 $117\pm 3$ $850\pm 19$ 382 0.009 Halo5 $145\pm 13$ $1676\pm 38$ 309 0.042
@@
Figure 14: The representative surface brightness profile of Halo4 is modeled using a two-component approach. The actual density profile is shown as black-filled circles at binned radii. The inner component is depicted in red, while the outer component, characterized by a scale radius of $r_{\rm s}=850$ pc, is illustrated in blue. The transition between the inner and outer components occurs at a radius of $r_{\rm trans}=382$ pc.
@@
In the top panels of Fig. 13, we illustrate the spatial distribution of our UFD analogs, with star particles from different progenitor halos represented in distinct colors overlaid on the surface brightness maps. The background stars used for fitting are excluded from this plot. Among the simulated UFD analogs, Halo1 exhibits the most compact features, with a half-light radius of $r_{\rm h,\ fit}\approx 87$ pc and an ellipticity of $\epsilon_{\rm fit}=0.34$, indicating a spherical distribution of stellar particles. Conversely, stars in Halo5 are the most extended, with $r_{\rm h,\ fit}\approx 754$ pc, and tend to deviate from a spherical shape, displaying a more elongated distribution with an ellipticity of $\epsilon_{\rm fit}=0.57$. Note that Halo6, which has a halo mass similar to that of Halo5, is excluded from the size relation analysis to avoid redundancy.
The bottom panels of Fig. 13 display the surface brightness profile as a function of the elliptical radius from the center of the simulated UFD analogs. The red circles (including background stars) and grey circles (excluding background stars) represent the surface brightness directly computed from stars in each UFD analog. The solid red line depicts the exponential fit derived using maximum likelihood estimation, including background stars, while the grey line represents the fit that excludes background stars. We verify that accurately estimating the spatial structures of simulated UFD analogs using observational methodologies necessitates the inclusion of background stars. Due to the presence of extended stars at large radii (indicated by grey circles), the fit without background stars (grey line) fails to capture the extended structure. This suggests that a different density profile, one that simultaneously accounts for both the inner, compact structure and the outer, extended structure, would be more appropriate (e.g., Kang & Ricotti, 2019; Ricotti et al., 2022; Jensen et al., 2024).
@@ 3.3.3 Size-luminosity relation
We compare the half-light radii of our simulated UFD analogs with both observational data and other theoretical studies in Fig. 17. The observational results include an updated catalog of dwarf galaxies from McConnachie (2012) and Simon (2019), as well as measurements of UFDs reported by Richstein et al. (2022, 2024). While Richstein et al. (2022, 2024) use azimuthally averaged half-light radii, we adopt the elliptical half-light radius, denoted as $a_{\rm h}$ in their work, to maintain consistency with other works using this definition. The simulated galaxies are represented in different colors (Jeon et al., 2017, 2021b; Wheeler et al., 2019; Applebaum et al., 2021; Prgomet et al., 2022; Sanati et al., 2023). For the V-band luminosity, we uniformly apply a mass-to-light ratio of 2 to our results and the other simulated galaxies.
@@
As illustrated in Fig. 17, we estimate the half-light radius of each UFD analog using two different methods: the direct method ($r_{\rm h,\ dir}$, shown as circles) and the two-component fitting method ($r_{\rm h}$, shown as crosses). While the $r_{\rm h}$ values are consistently smaller than the $r_{\rm h,\ dir}$ values for all UFD analogs, the former closely match the observational data. This consistency suggests that the large half-light radii reported in other simulations (e.g., Wheeler et al., 2019; Sanati et al., 2023) could be reduced by applying the two-component fitting method if extended structures are present.
Although our derived values using the two-component approach successfully match the observational data at the upper range, the average values remain larger ($r_{h}\sim$ a few $100\ \rm pc$) compared to the typical observed values ($r_{h}\lesssim 100\ \rm pc$). These larger sizes in the UFD regime predicted by theoretical work imply that observed UFDs may possess more extended yet unresolved structures. This possibility is further supported by recent observational studies revealing the presence of extended structures in UFDs (e.g., Tau et al., 2024; Jensen et al., 2024).
In particular, our simulations struggle to reproduce the highly compact sizes with $r_{\rm h}\lesssim 50$ pc. This discrepancy is particularly notable for UFDs with $M_{\star}\sim 10^{3}~{}\mbox{M}_{\odot}$, where observational studies consistently suggest sizes around $r_{\rm h}\sim 30$ pc. To explain the absence of compact sizes in the simulations, numerical heating has been suggested as a possible cause. This effect, which arises from the difference in mass resolution between dark matter and star particles, could contribute to the larger stellar sizes (e.g., Revaz & Jablonka, 2018; Binney & Knebe, 2002; Wilkinson et al., 2023).
@@
Specifically, Ludlow et al. (2019b) showed that the heavier DM particles tend to sink to the center of simulated galaxies, causing the less massive stellar particles to gradually diffuse outward. This process results in a systematic increase in the half-light radius over time until it reaches the convergence radius, $r_{\rm conv}\approx 0.055\times l(z)$ (Ludlow et al., 2019a), where $l(z)$ represents the mean interparticle spacing, defined as $l=L_{\rm box,\ phy}/N_{\rm part}^{1/3}$. Here, $L_{\rm box,\ phy}$ is the physical size of the simulation box, and $N_{\rm part}$ is the number of particles. In our study, the zoom-in region, with a resolution of a $2048^{3}$ grid, yields a convergence radius of $r_{\rm covn}\approx 59$ pc, including baryonic particles. Given that our estimated $r_{\rm h}$ values are larger than this limit, the impact of numerical heating may not be significant in our work. Furthermore, Wilkinson et al. (2023) show that the disk height in a MW-like galaxy can artificially increase due to spurious dynamical heating when the halo is resolved with fewer than $10^{6}$ DM particles. In this regard, our UFD analogs are unlikely to be affected by such spurious dynamical heating, as all halos except Halo1 are resolved with more than $10^{6}$ DM particles.
An additional point worth considering is the influence of collisional dynamics. As simulations of galaxy evolution achieve increasingly higher resolution, the resulting decrease in interstellar distances naturally raises the question of whether the collisionless treatment of stars remains valid. To date, collisional effects have primarily been studied in the context of star cluster formation. For example, Lahén et al. (2025) modeled cluster formation in an isolated disc galaxy using collisional dynamics and showed that this approach increases the effective radius of clusters by up to a factor of 10, bringing their sizes into better agreement with observations. However, due to the sub-parsec inter-particle distances involved in such clusters, it remains uncertain how collisional effects might impact stellar distributions in UFDs.
@@
We also mention that Jeon et al. (2021a, b), where a simulation setup similar to ours ($m_{\rm DM}=500~{}\mbox{M}_{\odot}$ and $m_{\rm gas}=63~{}\mbox{M}_{\odot}$), demonstrated the existence of simulated UFDs with smaller sizes ($r_{\rm h}<50$ pc). In these studies, numerical heating caused by differing mass ratios between DM and star particles was not a concern, as each star particle was modeled as a SSP with an initial mass of $m_{\star}\approx 500~{}\mbox{M}_{\odot}$, formed by accreting gas particles, giving rise to the same mass as a gas particle. However, the total stellar masses in these UFD analogs consisted of fewer than 10 particles for a stellar mass of $M_{\star}\lesssim 2\times 10^{3}~{}\mbox{M}_{\odot}$. This small number of particles representing the total stellar mass may contribute to the lower $r_{\rm h}$ values.
@@
Figure 18: Stellar metallicity as a function of 2D elliptical radius from the center at $z=0$. Here, the metallicity gradient is not observed in any simulated UFDs; rather, these galaxies exhibit a wide range of metallicity distributions, reflecting the diverse origins of stars from different progenitor halos. Specifically, in Halo5, stars originating from prog5 (indicated in purple) are found dispersed across vast regions extending from 1 to 10 kpc.